# Collaborative Aquatic Positioning system Utilising Multi-beam Sonar and Depth Sensors

### 15 Mar 2024 v1

@@ Abstract
Accurate positioning of underwater robots in confined environments is crucial for inspection and mapping tasks and is also a prerequisite for autonomous operations. Presently, there are no positioning systems available that are suited for real-world use in confined underwater environments, unconstrained by environmental lighting and water turbidity levels and have sufficient accuracy for reliable and repeatable navigation. This shortage presents a significant barrier to enhancing the capabilities of ROVs in such scenarios. This paper introduces an innovative positioning system for ROVs operating in confined, cluttered underwater settings, achieved through the collaboration of an omnidirectional surface vehicle and an ROV. A formulation is proposed and evaluated in the simulation against ground truth. The experimental results from the simulation form a proof of principle of the proposed system and also demonstrate its deployability. Unlike many previous approaches, the system does not rely on fixed infrastructure or tracking of features in the environment and can cover large enclosed areas without additional equipment.
@@ III-A Notations and coordinate frames
Figure 1: Overview of the proposed confined aquatic positioning (CAP-SD) system. An omnidirectional ASV - MallARD, with integrated LiDAR, IMU, and an image sonar tracks a submersible ROV. The ROV is also fitted with a pressure sensor for depth measurement.
Figure 1 presents the coordinate system and reference frames used in this work. The full positioning system is composed of the following co-ordinate frames: world frame $\mathcal{F}_{W}$, ASV baselink frame $\mathcal{F}_{B}$, multi-beam sonar $\mathcal{F}_{S}$ and underwater robot frame $\mathcal{F}_{R}$. The world frame $\mathcal{F}_{W}$ origin is assigned to a corner of the testing tank. $\mathcal{F}_{B}$ is the geometric centre of the ASV. The origin of the sonar frame $\mathcal{F}_{S}$ is located at the optical center of the transmitter and has a fixed transform from $\mathcal{F}_{B}$.
@@
III-B CAP-SD system formulation
The proposed CAP-SD system initially performs self-localisation of the surface robot based on the SLAM (Simultaneous Localisation and Mapping) and the integrated IMU (Inertial Measurement Unit), to ascertain the pose of the surface robot relative to the world-fixed frame. And then, by employing the multi-beam sonar mounted on the surface robot, the system localises the underwater robot, thereby determining its pose relative to the surface robot frame. This process enables the determination of the underwater robot’s pose within the world-fixed frame. However, the pose of underwater robot can not be defined by this pipeline. Multi-beam sonar systems generate 2-D visualisations of the 3-D environment, where each pixel delineates an arc originating from the center of the multi-beam sonar’s transducer. In this scenario, the pixels have a direct correlation with both the azimuth angle ($\theta$) and the distance (R) of the arc from its origin, but the aggregation of spatial returns along these arcs precludes the retrieval of elevation angles ($\phi$), as shown in Figure 2A. In this scenario, the CAP-SD system utilizes depth sensors mounted on the underwater robot to define a plane (depth plane). The intersection of this plane with the arc is used to determine the pose of the underwater robot.
First, the pose of the sonar with respect to world-fixed frame ($\mathcal{F}_{W}$), which also serves as the arc’s center, can be expressed by (detailed in Section III-D):
@@ III-B CAP-SD system formulation
Second, following the principles of multi-beam sonar imaging, it is possible to determine the distance from the sonar to the underwater robot R. Additionally, this method yields the angle $\theta$, representing the orientation of the cutting-sphere plane relative to the $x-z$ plane around the sonar’s $y$ axis in the sonar frame (detailed in Section III-C). The sphere is defined as:
@@
where $x_{S}^{W}$, $y_{S}^{W}$ and $z_{S}^{W}$ is the position of multi-beam sonar. The cutting-sphere plane containing the arc can be represented as (gray plane depicted in Figure 2B):
@@
where $\vec{n}_{3\times 1}=[A,B,C]^{\top}$ is the normal vector to the cutting-sphere plane, defined as:
$\left[\begin{array}[]{c}\vec{n}_{3\times 1}\\ 1\end{array}\right]=(H^{W}_{B}H^{B}_{S})^{\prime}\left[\begin{array}[]{c}\sin(\theta)\\ -\cos(\theta)\\ 0\\ 1\end{array}\right].$ (5)
To determine the position of the underwater robot, it is necessary to utilize a depth sensor that can abstractly define a plane:
@@
By intersecting the arc (gray circle in Figure 2C), derived from solving Equation 3 and 4, with the plane defined by the depth sensor, two points can be obtained. However, the target underwater robot to be located must appear within the FOV of the image sonar (sonar’s positive x region as shown orange arc in Figure 2C), When transformed into the world-fixed frame, the $x^{W}_{R}$ and $y^{W}_{R}$ must satisfy the following condition:
@@
Figure 2: Abstract frames of the CAP-SD system. (A) Underwater side-view of CAP-SD system. (B) Using the cutting-sphere plane and sphere to define the arc where the underwater robot may be located. (C) Possible locations of underwater robot can be reduced to two points ($P_{1}$ and $P_{2}$) by generating another plane using the depth sensor. The position of underwater robot ($P_{2}$) must be in the sonars positive x region (orange arc in (B) and (C)).
III-C Neural network (NN) for image sonar object detection
@@
III-D Calculating ASV rotation relative to the world frame $\mathbf{R}^{W}_{B}$
@@
The position and orientation data $[x^{W}_{B},y^{W}_{B},\psi^{W}_{B}]^{\top}$ provided by SLAM is notably accurate, as verified by our prior research [27]. The fusion of SLAM and IMU data results in Euler angles $[\psi^{W}_{B},\theta^{W}_{B},\phi^{W}_{B}]^{\top}$, with pitch ($\theta$) and roll ($\phi$) derived from the tilting EKF. Consequently, the rotation matrix $\mathbf{R}^{W}_{B}$ is formulated using the conventional Z-Y-X Euler angles based on the fused orientation data.
@@ IV-B Simulation Experiment Scenario
Figure 5: Simulation environment scenario and robots. (A) Overview of the simulated nuclear fuel pool (B) Sideoview of the nuclear fuel pool (B) Autonomous surface robot - MallARD (D) MallARD’s 3D model in Unity (E) Commercial underwater robot - BlueROV2 (F) BlueROV2 3D model in Unity.
@@ IV-D Discussion
Figure 6B shows the 3D trajectory of the CAP-SD alongside the ground truth. Overall, within this simulation environment, CAP-SD is capable of accurately localising the underwater robot. Table 3 presents the RMS error of the proposed CAP-SD methods for all five datasets. Due to the depth/pressure plugin used in this simulation not actually sensing water pressure but merely adding noise to a known absolute depth, the results do not include a comparison of the RMSE along the $z$-axis. Therefore, as detailed in Section III, the accuracy of localisation largely depends on whether the YOLOv5 model can accurately determine the bounding box (the pixel coordinates of the robot on the sonar image frame).
As indicated by the Table 3, the Euclidean RMSE remains below 200 mm for each trajectory type. In datasets 2 and 3, the values are relatively higher, approaching 200 mm. This is attributed to the BlueROV2 passing through specific locations during the bouncing and lawnmower trajectories, where, in the sonar images, the spatial overlap between the robot and other objects in the underwater environment occurs. Under these circumstances, YOLOv5 fails to provide accurate pixel coordinates for the robot, as shown in Fig 7. In Fig 7A, the robot is not significantly affected by the environment, allowing YOLOv5 to clearly detect the BlueROV2 in the sonar images. However, in certain areas, as shown in Fig 7B and C, this is not the case. The green bounding box indicates the expected range, while the red represents the results from YOLOv5.
Figure 7: Sonar image view with YOLOv5 object detection. (A) Accurate robot detection by YOLOv5 (B) and (C) Inaccurate robot detection by YOLOv5.
@@
TABLE I: RMS errors relative to the ground truth measurement of CAP-SD for all 5 datasets  [^3] All datasets are available in: https://github.com/Xue1iang/IROS2024_CAP-SD.

### 18 Mar 2024 v2

@@ Abstract
Accurate positioning of remotely operated underwater vehicles (ROVs) in confined environments is crucial for inspection and mapping tasks and is also a prerequisite for autonomous operations. Presently, there are no positioning systems available that are suited for real-world use in confined underwater environments, unconstrained by environmental lighting and water turbidity levels and have sufficient accuracy for long-term, reliable and repeatable navigation. This shortage presents a significant barrier to enhancing the capabilities of ROVs in such scenarios. This paper introduces an innovative positioning system for ROVs operating in confined, cluttered underwater settings, achieved through the collaboration of an omnidirectional surface vehicle and an ROV. A formulation is proposed and evaluated in the simulation against ground truth. The experimental results from the simulation form a proof of principle of the proposed system and also demonstrate its deployability. Unlike many previous approaches, the system does not rely on fixed infrastructure or tracking of features in the environment and can cover large enclosed areas without additional equipment.
@@ III-A Notations and coordinate frames
Figure 1: Overview of the components and coordinate frames of the confined aquatic positioning (CAP-SD) system. An omnidirectional ASV - MallARD, with integrated LiDAR, IMU, and an image sonar tracks a submersible ROV. The ROV is also fitted with a pressure sensor for depth measurement.
Figure 1 presents the coordinate system and reference frames used in this work. The full positioning system is composed of the following co-ordinate frames: the inertial world frame $\mathcal{F}_{W}$, ASV baselink frame $\mathcal{F}_{B}$, multi-beam sonar $\mathcal{F}_{S}$ and ROV frame $\mathcal{F}_{R}$. The world frame $\mathcal{F}_{W}$ origin is assigned to a corner of the testing tank. $\mathcal{F}_{B}$ is the geometric centre of the ASV. The origin of the sonar frame $\mathcal{F}_{S}$ is located at the optical center of the transmitter and has a static transform from $\mathcal{F}_{B}$.
@@
III-B CAP-SD formulation overview
The full transform chain from the world frame $\mathcal{F}_{W}$ to the ROV frame $\mathcal{F}_{R}$:
$\mathbf{H}_{S}^{W}=\mathbf{H}_{B}^{W}\mathbf{H}_{S}^{B}\mathbf{H}^{S}_{R},$ (2)
can not be used to obtain $\mathbf{P}^{W}_{R}=[x^{W}_{R},y^{W}_{R},z^{W}_{R}]^{\top}$ directly because $\mathbf{P}^{S}_{R}$ cannot be defined. This is because the multi-beam sonar generates 2-D visualisations of the 3-D environment, where each pixel in the image represents an arc in 3D space, originating from the center of the multi-beam sonar.
In this scenario, each pixel has a direct correlation with both the azimuth angle ($\theta$) and the distance (R) of the arc from its origin, but the aggregation of spatial returns along these arcs precludes the retrieval of elevation angles ($\phi$), as shown in Figure 2A. Therefore, even if the location of the ROV is successfully identified in the sonar’s scan, the translation $\mathbf{P}^{S}_{R}=[x^{S}_{R},y^{S}_{R},z^{S}_{R}]^{\top}$ can not be defined and the transform chain is incomplete. To resolve this problem, the CAP-SD system uses the available elements of the transform chain and the pixel location of the ROV in the sonars image to generate an arc of possible positions of the ROV in the world frame $\mathcal{F}_{W}$. A pressure sensor on the ROV is used to generate a horizontal plane (depth plane) of possible robot locations in $\mathcal{F}_{W}$. By finding the intersection of the arc and the horizontal plane, a single location for the ROV in  $\mathcal{F}_{W}$ can be ascertained.
III-C CAP-SD formulation details
First, the pose of the sonar with respect to world-fixed frame ($\mathcal{F}_{W}$), which also serves as the arc’s center, can be expressed by:
@@ III-C CAP-SD formulation details
where $\mathbf{H}_{S}^{B}$ is a static homogenous transformation from ASV’s baselink frame to sonar frame and $\mathbf{H}_{B}^{W}$ is detailed in Section III-E.
Second, the pixel location of the ROV in the sonar scan is identified using an object detection neural network. The pixel location in the sonar image is used to determine the distance from the sonar to the ROV R and the angle $\theta$, as detailed in Section III-D..
Third, the arc, represented by a circle in 3D space, is defined. This involves first defining a sphere of possible locations and then defining a sphere-cutting plane, where the intersection between the sphere and the sphere-cutting plane defines the circle. The centre of the sphere is the origin of $\mathcal{F}_{S}$ and it’s radius is $R$. The sphere is defined as:
@@
where $[x_{S}^{W}y_{S}^{W}z_{S}^{W}]^{\top}$ is the position of multi-beam sonar. The sphere-cutting plane can be defined by taking the plane of the $x-z$ axes of $\mathcal{F}_{S}$ and rotating it by the angle $\theta$ about the $y$ axis of $\mathcal{F}_{S}$ (gray plane depicted in Figure 2B). The sphere-cutting plane containing the arc can be represented as:
@@
where $\vec{\mathbf{n}}_{3\times 1}=[A,B,C]^{\top}$ is the normal vector to the cutting-sphere plane, defined as:
$\left[\begin{array}[]{c}\vec{\mathbf{n}}_{3\times 1}\\ 1\end{array}\right]=(H^{W}_{B}H^{B}_{S})^{\prime}\left[\begin{array}[]{c}\sin(\theta)\\ -\cos(\theta)\\ 0\\ 1\end{array}\right].$ (6)
Fourth, the depth plane in $\mathcal{F}_{W}$ must be defined. This is achieved by attaching a pressure sensor to the ROV and calibrating it to measure depth below the surface. The depth plane, which is normal to the $\mathcal{F}_{W}$ $z$-axis, can be defined as:
@@
where $z_{depth}$ is the depth measured by the pressure sensor.
Fifth, by intersecting the sphere, sphere-cutting plane, and the depth plane, the solution domain can be reduced to two points in $\mathcal{F}_{W}$. This is achieve by solving equations (4), (5) and (7) for $x,y$ and $z$.
The final step is to ascertain which of the two points is the correct one. This is relatively simple as the correct solution lies in the positive $x$ region in $\mathcal{F}_{S}$, as shown orange arc in Figure 2C). When transformed into the world-fixed frame, the $x^{W}_{R}$ and $y^{W}_{R}$ must satisfy the following condition:
@@
Figure 2: Abstract frames of the CAP-SD system. (A) Underwater side-view of CAP-SD system. (B) Using the cutting-sphere plane and sphere to define the arc where the ROV may be located. (C) Possible locations of ROV can be reduced to two points ($P_{1}$ and $P_{2}$) by generating another plane using the depth sensor. The position of ROV ($P_{2}$) must be in the sonars positive x region (orange arc in (B) and (C)).
III-D Neural network (NN) for image sonar object detection
@@
III-E Calculating ASV rotation relative to the world frame $\mathbf{R}^{W}_{B}$
@@
The position and orientation data $[x^{W}_{B},y^{W}_{B},\psi^{W}_{B}]^{\top}$ provided by SLAM (Hector SLAM [27] was deployed on the ASV in the simulation) is notably accurate, as verified by our prior research [28]. The fusion of SLAM and IMU data results in Euler angles $[\psi^{W}_{B},\theta^{W}_{B},\phi^{W}_{B}]^{\top}$, with pitch ($\theta$) and roll ($\phi$) derived from the tilting EKF. Consequently, the rotation matrix $\mathbf{R}^{W}_{B}$ is formulated using the conventional Z-Y-X Euler angles based on the fused orientation data.
@@ IV-B Simulation Experiment Scenario
Figure 5: Simulation environment scenario and robots. (A) Overview of the simulated nuclear fuel pool  [^3] The CAD model of nuclear fuel pool used in this simulation is derived from the underground section (Bomb site B https://www.youtube.com/watch?v=wzbfPDgN5lo) of a map from the online game Counter-Strike 2, Nuke.: https://steamcommunity.com/sharedfiles/filedetails/?id=2324376136. (B) Sideoview of the nuclear fuel pool (B) Autonomous surface robot - MallARD (D) MallARD’s 3D model in Unity (E) Commercial underwater robot - BlueROV2 (F) BlueROV2 3D model in Unity.
@@ IV-D Discussion
Figure 6B shows the 3D trajectory of the CAP-SD alongside the ground truth. Overall, within this simulation environment, CAP-SD is capable of accurately localising the ROV. Table 5 presents the RMS error of the proposed CAP-SD methods for all five datasets. Due to the depth/pressure plugin used in this simulation not actually sensing water pressure but merely adding noise to a known absolute depth, the results do not include a comparison of the RMSE along the $z$-axis. Therefore, as detailed in Section III, the accuracy of localisation largely depends on whether the YOLOv5 model can accurately determine the bounding box (the pixel coordinates of the robot on the sonar image frame).
As indicated by the Table 5, the Euclidean RMSE remains below 200 mm for each trajectory type. In datasets 2 and 3, the values are relatively higher, approaching 200 mm. This is attributed to the BlueROV2 passing through specific locations during the bouncing and lawnmower trajectories, where, in the sonar images, the spatial overlap between the robot and other objects in the underwater environment occurs. Under these circumstances, YOLOv5 fails to provide accurate pixel coordinates for the robot, as shown in Fig 7. In Fig 7A, the robot is not significantly affected by the environment, allowing YOLOv5 to clearly detect the BlueROV2 in the sonar images. However, in certain areas, as shown in Fig 7B, this is not the case. The green bounding box indicates the expected range, while the red represents the results from YOLOv5. This is also a challenge for multi-beam sonar: signal-to-noise ratio (SNR) and resolution. Images from multi-beam sonar frequently exhibit artifacts and possess a lower resolution when contrasted with images captured by cameras [33].
Figure 7: Sonar image view with YOLOv5 object detection. (A) Accurate robot detection by YOLOv5. (B) Due to YOLO v5’s inability to provide precise bounding boxes, errors occur in localisation. Green represents the true position, while red indicates the detection results of YOLOv5.
@@
TABLE I: RMS errors relative to the ground truth measurement of CAP-SD for all 5 datasets  [^5] The source code and details about the rest dataset are available in: https://github.com/Xue1iang/IROS2024_CAP-SD.